OpenStudy (anonymous):

What is the solution to the system of equations? x=3y+5 x-4y=3 (_,_)

OpenStudy (anonymous):

put the \(3y+5\) in the first equation where you see \(x\) in the second equation and start with \[3y+5-4y=3\] then solve that one for \(y\)

OpenStudy (anonymous):

so subtract y from both sides

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

what is y

OpenStudy (anonymous):

HELp

OpenStudy (anonymous):

:)

OpenStudy (javk):

So you know what x is equal to so like @satellite73 said substitue that into your second eqn \[3y+5-4y=3\] take all the nos to one side n expressions to the other \[3y-4y=3-5\] \[-y=-2\] therefore \[y=2\] that you what y is solve for x \[x=3y+5\] \[x=3(2)+5\] and solve

OpenStudy (anonymous):

x=11